OpenStudy (anonymous):

PLEASE HELP, DUE TODAY: Find the cosine of the angle between the curves cos t, sin(t)/4, sin tand cos t, sin t, sin(2t)where they intersect. The possible answers are : 7√5√17/85 or −9√5√17/85. Please show all of your work. Thanks

OpenStudy (anonymous):

I won't do it for you, but try taking the derivative in the intersection point, and remember that it is the tangent of the angle between the horizontal and the tangent line

OpenStudy (anonymous):

Okay, how do I find the intersection point? I was told to do this: cos(t)=cos(u) sin(t)/4=sin(u) sin(t)=sin(2u) So I solve for u in one of the equations, and substitute it into one of the other equations and that is supposed to get me a point of intersection...but it just doesnt seem right to say u=arccos(cos(t)).. I'm really confused... can you explain a little bit more?

OpenStudy (anonymous):

You are right, that is really strange and I believe it is wrong. Can you put more clearly the functions of the heading, because I'm confused with what you wrote.

OpenStudy (anonymous):

Sure. \[<\cos(t), \frac{ -\sin(t) }{ 4 }, \sin(t)> \] and \[<\cos(t), \sin(t), \sin(2t)>\]. Those are the two curves. So the point of intersection comes from setting x_1=x_2, y_1=y_2, and z_1=z_2, but like I said before, it doesnt look right...

OpenStudy (anonymous):

Aaah ok, now I understand. That is right, you need to solve those equations now. Was that your doubt, or you need help with this part?

OpenStudy (anonymous):

Yeah, that's the part I need help in :/

OpenStudy (anonymous):

Well cos(t) is always the same, what about the third equation, when is sin(t)=sin(2t)? What can you do here?

OpenStudy (anonymous):

\[\frac{ \arcsin(\sin(t)) }{ 2 }=u\]. Is that any better?

OpenStudy (anonymous):

try logic, think about the trigonometric circle, this way it gets more complicated

OpenStudy (anonymous):

t=pi?

OpenStudy (anonymous):

Yes, that is a particular solution, and maybe the only one you need, but te general solution is t=n pi, where n is any integer, so 2pi is also an important solution. Ok now go to the second equation. Remember that for both curves intersect, all 3 equations must be satisfied, so the right solution for the second one must also be a solution for the third, but try to do the general solution first.

OpenStudy (anonymous):

Im confused again :/ what do i do with t=n*pi now? I'm sorry, I've just been stuck on this problem for about 3 hours now

OpenStudy (anonymous):

Dont worry about n pi, solve the second one, then you come back to n pi. In mathematics never try to think of how are you going to use that latter if you know that you did what you had to do. You need to solve those equations, that is obvious and right, so don't worry about later. Only do that if you think you already screwed up.

OpenStudy (anonymous):

Well, another solution would be pi, wouldnt it?

OpenStudy (turingtest):

how would you guys propose to when y is the same?

OpenStudy (turingtest):

to find*

OpenStudy (anonymous):

I don't understand the question, can you rephrase it?

OpenStudy (turingtest):

actually nevermind I was being slow :/ it's easy

OpenStudy (anonymous):

Can you help me out?

OpenStudy (turingtest):

I think so, yeah no interval is given for t?

OpenStudy (anonymous):

Nope

OpenStudy (turingtest):

well I guess maybe the shape is periodic so the angles repeat you guys seem to be on the right track cos t=cos t -> t is all reals sin t=1/4sin t ->t=0 sin t=sin(2t) ->t=n*pi, n=0,1,2...

OpenStudy (turingtest):

only one value of t fits all three solutions, so...?

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

yeah, looks like it

OpenStudy (turingtest):

then do what @ivanmlerner said and get the derivatives (i.e. tangent vectors) at that point

OpenStudy (anonymous):

How do i get the derivatives if i only have t=0?

OpenStudy (turingtest):

take the derivatives first

OpenStudy (turingtest):

...before plugging in the numbers that is the tangent vector function

OpenStudy (anonymous):

Wait, for the second solution t is not only 0, it can be any n*pi also but since the angles repeat that does not make any difference

OpenStudy (turingtest):

ah, you are right, but the choice of 0 or n*pi should give the same answer, no?

OpenStudy (anonymous):

So i could really use either, because there are two possible answers?

OpenStudy (turingtest):

0 or n*pi are the only solutions for the y and z equations the sines and cosines will have the same values, and since we only have sint and cost in both our position function and tangent function I argue that it makes no difference whether we use t=0 or t=n*pi

OpenStudy (turingtest):

though I have been known to be wrong in my day, my theory is that this shape is periodic, and we will get the same answer either way let's try and see if we can get one of your choices...

OpenStudy (anonymous):

Okay, so, let's try 0. I plug 0 in into \[<\cos(t), \sin(t), \sin(2t)\] and get \[<1, 0, 0>\] and that would be hwere the curves meet, right?

OpenStudy (anonymous):

And now, the derivatives are \[<-\sin(t), -\frac{ \cos(t) }{ 4 }, \cos(t)> and <-\sin(t), \cos(t), 2\cos(2t)>\] yes?

OpenStudy (anonymous):

Oh my god that is so cool, I graphed it and maybe it would be better to check 0 and pi.

OpenStudy (turingtest):

@kandsy yes

OpenStudy (turingtest):

the y component in the first function is negative? I missed that

OpenStudy (anonymous):

yes, because -cos is the deriv of -sin

OpenStudy (turingtest):

right, but I didn't see the -sin in your post, I think there is a typo

OpenStudy (anonymous):

Right, i see that now, my original post was the typo, it does in fact have a negative.

OpenStudy (turingtest):

My answer is close to your options but I'm ot sure where the 85 is coming from anyway let's continue, you're doing fine...

OpenStudy (turingtest):

oh I see it's just rationalized yes I have one answer...

OpenStudy (turingtest):

I have one issue which is that I am not getting the negative sign

OpenStudy (turingtest):

I hate missing negative signs :S

OpenStudy (turingtest):

oh wait, now I got it okay, we're good :)

OpenStudy (anonymous):

so now i plug 0 into the first curve and get \[<-\sin(0),\frac{ −\cos(0)}{ 4 },\cos(0)>\] right? And so I get: <0, -1/4, 1> and that is one of the points of intersection. Now i plug pi into the other curve, right?

OpenStudy (turingtest):

the best way to do it in my opinion is to do the dot product of the two vectors without pluggin in the numbers until the end

OpenStudy (turingtest):

so let\[\vec r_1'=\langle-\sin t,-\frac14\cos t,\cos t\rangle\]\[\vec r_2=\langle-\sin t,\cos t,2\cos(2t)\cos t\rangle\]

OpenStudy (turingtest):

now apply the formula\[\vec r_1\cdot\vec r_2=\|\vec r_1\|\|\vec r_2\|\cos\theta_{12}\]

OpenStudy (turingtest):

stupid LaTeX :/

OpenStudy (turingtest):

I mean\[\vec r_1'\cdot\vec r_2'\]then plug in t=0 excuse me, I am sorry to confuse you

OpenStudy (turingtest):

that should make more sense though, as those are the tangent vectors

OpenStudy (anonymous):

Thank you VERY much :) It's made a lot of sense. You're fine. But, what youre telling me is, that i only have to plug 0 in, in order to get one of the answers? Because, I only need one.

OpenStudy (turingtest):

then do the same for\[\|\vec r_1'\|\]and\[\|\vec r_2'\|\]

OpenStudy (turingtest):

yes you only need to plug in t=0, I already checked

OpenStudy (turingtest):

then once you got all that business, finally apply\[\vec r_1'\cdot\vec r_2'=\|\vec r_1'\|\|\vec r_2'\|\cos\theta_{12}\]solve for \(\cos\theta_{12}\) and you're done :)

OpenStudy (anonymous):

Okay, great, thanks. My problem has been solved :)

OpenStudy (turingtest):

you are welcome :D happy to help!